# Bike-Bench: A Bicycle Design Benchmark for Generative Models with Objectives and Constraints

### v1

@@ A.4 Summary of Constraints and Objectives
We describe Bike-Bench’s constraints and objectives below. A summary of all evaluation functions is included in Table 4, and a text description of each is included as follows:
  • Usability Score: [Objective] The predicted ‘usability,’ as rated by a human, with 0 being the most usable and 1 being the least usable. Predicted by a regression model trained on human-collected ratings.
  • Drag Force: [Objective] The predicted drag force in N incurred by the cyclist in a 10 m/s headwind, as predicted by a regression model trained on computational fluid dynamics simulation data.
  • Knee Angle Error: [Objective] The difference between the minimum knee angle of the cyclist and the optimal reference range. May include a penalty term if the rider’s geometry is completely incompatible with the bike.
  • Hip Angle Error: [Objective] The difference between the torso-to-upper-leg angle of the cyclist and the optimal reference range. May include a penalty term if the rider’s geometry is completely incompatible with the bike.
  • Arm Angle Error: [Objective] The difference between the torso-to-arm angle of the cyclist and the optimal reference range. May include a penalty term if the rider’s geometry is completely incompatible with the bike.
  • Cosine Distance to Embedding: [Objective] The cosine distance in the CLIP embedding space between the rendered bike image and the target text or image embedding.
  • Mass: [Objective] The mass in kg of the bike frame, as predicted by a regression model trained on finite element analysis data.
  • Planar Compliance: [Objective] A composite planar compliance score for the bike frame, as predicted by a regression model trained on finite element analysis data.
  • Transverse Compliance: [Objective] A transverse compliance score for the bike frame, as predicted by a regression model trained on finite element analysis data.
  • Eccentric Compliance: [Objective] A composite eccentric compliance score for the bike frame, as predicted by a regression model trained on finite element analysis data.
  • Planar Safety Factor: [Constraint] Constraint quantified as 1.5 minus the safety factor under planar loading, as predicted by a regression model trained on finite element analysis data.
  • Eccentric Safety Factor: [Constraint] Constraint quantified as 1.5 minus the safety factor under eccentric loading, as predicted by a regression model trained on finite element analysis data.
  • Saddle height too small: [Constraint] Constraint indicating that the saddle height collides with the top of the seat tube.
  • Seat post too short: [Constraint] Constraint indicating that the seat post doesn’t reach the seat tube given the prescribed saddle height.
  • Head tube lower extension too great: [Constraint] Constraint indicating that the down tube doesn’t properly intersect with the head tube.
  • Head tube length too great: [Constraint] Constraint indicating that the head tube is so short that the top tube and down tube intersect.
  • Certain parameters must be positive: [Constraint] Constraint indicating that at least one parameter that should be strictly positive is negative.
  • Chain stay should be greater than wheel radius: [Constraint] Constraint indicating that the chain stay length is smaller than the wheel radius, creating a collision.
  • Chain stay should be greater than BB: [Constraint] Constraint indicating that the vertical drop from the rear axle to bottom bracket is greater than the chain stay length, creating an impossibility.
  • Seat stay should be greater than wheel radius: [Constraint] Constraint indicating that the seat stay length is smaller than the wheel radius, creating a collision.
  • Down tube must reach head tube: [Constraint] Constraint indicating that the down tube is too short to reach the head tube.
  • The pedal shouldn’t intersect the front wheel: [Constraint] Constraint indicating that the front wheel would intersect the pedal in its forward position, causing a collision when turning.
  • The crank shouldn’t hit the ground when it is in its lower position: [Constraint] Constraint indicating that the crank hits the ground during its rotation.
  • RGB value should be less than 255: [Constraint] Constraint indicating that frame RGB values were set at higher than 255.
  • Predicted Frame Validity: [Constraint] Constraint indicating some abstract issue with the frame, as predicted by a classification model trained to identify CAD models that failed to regenerate.
A.5 Summary of Design Variables
We include a text descriptions of all 70 parameters in the standard bike–bench design representation scheme. All lengths are measured in mm. All angles are measured in degrees. Parameter names are largely based on the BikeCAD internal XML labeling scheme.
The 70 variables are described as follows:
  • ‘CS textfield’: [Continuous] The length of the chain stay tubes.
  • ‘BB textfield’: [Continuous] Bottom bracket drop, measured as the vertical drop from the rear axle to the center of the bottom bracket. By convention, positive values imply the bottom bracket lies below the axle.
  • ‘Stack’: [Continuous] The vertical distance from the top of the head tube relative to the bottom bracket.
  • ‘Head angle’: [Continuous] The angle of the head tube with respect to horizontal, in degrees.
  • ‘Head tube length textfield’ [Continuous] The length of the head tube.
  • ‘Seat stay junction0’: [Continuous] The length along the seat tube from the top of the seat tube to the junction with the seat stays. By convention, this is measured to the center of the seat stays.
  • ‘Seat tube length’: [Continuous] The length of the seat stay tubes.
  • ‘Seat angle’: [Continuous] The angle of the seat tube with respect to horizontal.
  • ‘DT Length’: [Continuous] The length of the down tube.
  • ‘FORK0R’: [Continuous] Fork offset, measured as the perpendicular distance from the front axle to the head tube axis.
  • ‘BB diameter’: [Continuous] The diameter of the bottom bracket
  • ‘ttd’: [Continuous] Top tube outer diameter.
  • ‘csd’: [Continuous] Chain stay outer diameter.
  • ‘ssd’: [Continuous] Seat stay outer diameter.
  • ‘dtd’: [Continuous] Down tube outer diameter.
  • ‘Chain stay position on BB’: [Continuous] The distance along the length of the bottom bracket from its edge to the center of the chain stay tubes.
  • ‘SSTopZOFFSET’: [Continuous] The offset from center plane of the bike of the joints connecting the seat stays to the the seat tube.
  • ‘Head tube upper extension2’: [Continuous] The length from the top of the head tube to the junction with the top tube. By convention, this is measured to the center of the top tube.
  • ‘Seat tube extension2’: [Continuous] The length from the top of the seat tube to the junction with the top tube. By convention, this is measured to the center of the top tube.
  • ‘Head tube lower extension2’: [Continuous] The length from the bottom of the head tube to the junction with the down tube. By convention, this is measured to the center of the down tube.
  • ‘SEATSTAYbrdgshift’: [Continuous] The distance along the center plane of the bike from the seat stay and seat tube junction to the seat stay bridge, if present on the bike.
  • ‘CHAINSTAYbrdgshift’: [Continuous] The distance along the center plane of the bike from the outer rim of the bottom bracket to the chain stay bridge, if present on the bike.
  • ‘SEATSTAYbrdgdia1’: [Continuous] The diameter of the seat stay bridge, if present on the bike.
  • ‘CHAINSTAYbrdgdia1’: [Continuous] The diameter of the chain stay bridge, if present on the bike.
  • ‘SEATSTAYbrdgCheck’: [Boolean] A boolean value indicating whether the seat stay bridge is present on the bike.
  • ‘CHAINSTAYbrdgCheck’: [Boolean] A boolean value indicating whether the chain stay bridge is present on the bike.
  • ‘Dropout spacing’: [Continuous] The distance between the rear dropouts.
  • ‘Wall thickness Bottom Bracket’: [Continuous] The tube wall thickness of the bottom bracket.
  • ‘Wall thickness Top tube’: [Continuous] The tube wall thickness of the top tube.
  • ‘Wall thickness Head tube’: [Continuous] The tube wall thickness of the head tube.
  • ‘Wall thickness Down tube’: [Continuous] The tube wall thickness of the down tube.
  • ‘Wall thickness Chain stay’: [Continuous] The tube wall thickness of the chain stay.
  • ‘Wall thickness Seat stay’: [Continuous] The tube wall thickness of the seat stay.
  • ‘Wall thickness Seat tube’: [Continuous] The tube wall thickness of the seat tube.
  • ‘Wheel diameter front’: [Continuous] The outer diameter of the front wheel.
  • ‘RDBSD’: [Continuous] The difference between rear wheel outer diameter and bead seat diamater, roughly approximating the tire thickness.
  • ‘Wheel diameter rear’: [Continuous] The outer diameter of the rear wheel.
  • ‘FDBSD’: [Continuous] The difference between front wheel outer diameter and bead seat diamater, roughly approximating the tire thickness.
  • ‘Display AEROBARS’: [Boolean] A boolean value indicating whether the bike has aerobars.
  • ‘BB length’: [Continuous] The length of the bottom bracket.
  • ‘Head tube diameter’: [Continuous] Head tube outer diameter.
  • ‘Wheel cut’: [Continuous] The diameter of the cutout of seat tube for the rear wheel, if using an aerodynamic tube type.
  • ‘Seat tube diameter’: [Continuous] Seat tube outer diameter.
  • ‘bottle SEATTUBE0 show’: [Boolean] A boolean value indicating whether the bike has a bottle holder on the seat tube.
  • ‘bottle DOWNTUBE0 show’: [Boolean] A boolean value indicating whether the bike has a bottle holder on the down tube.
  • ‘Front Fender include’: [Boolean] A boolean value indicating whether the bike has a front fender.
  • ‘Rear Fender include’: [Boolean] A boolean value indicating whether the bike has a rear fender.
  • ‘BELTorCHAIN’: [Boolean] A boolean value indicating whether the bike has a chain (True) as opposed to a belt.
  • ‘Number of cogs’ [Integer] The number of cogs on the rear wheel.
  • ‘Number of chainrings’ [Integer] The number of chainrings attached to the crank.
  • ‘Display RACK’: [Boolean] A boolean value indicating whether the bike has a rack.
  • ‘FIRST color R_RGB’: [Continuous] The red component of the primary paint color of the bike.
  • ‘FIRST color G_RGB’: [Continuous] The green component of the primary paint color of the bike.
  • ‘FIRST color B_RGB’: [Continuous] The blue component of the primary paint color of the bike.
  • ‘SPOKES composite front’: [Integer] If applicable, the number of composite spokes in the front wheel minus two (a value of 1 is a trispoke wheel).
  • ‘SPOKES composite rear’: [Integer] If applicable, the number of composite spokes in the rear wheel minus two (a value of 1 is a trispoke wheel).
  • ‘SBLADEW front’: [Continuous] If applicable, the width of the front wheel composite spokes.
  • ‘SBLADEW rear’: [Continuous] If applicable, the width of the rear wheel composite spokes.
  • ‘Saddle length’: [Continuous] The length of the saddle.
  • ‘Saddle height’: [Continuous] The bertical distance from the saddle to the bottom bracket.
  • ‘Down tube diameter’: [Continuous] The diameter of the down tube.
  • ‘Seatpost LENGTH’: [Continuous] The length of the seat post.
  • ‘MATERIAL’: [Categorical] The material of the bike frame. Possible values are: ‘ALUMINIUM’, ‘CARBON’, ‘STEEL’, ‘TITANIUM’, ‘BAMBOO’, ‘OTHER’.
  • ‘Head tube type’: [Categorical] The style of head tube. Possible values are: ‘0’, ‘1’, ‘2’, ‘3’. 0 is aerodynamic, while 1 and 2 are standard round tubes with no distinction in this representation scheme. 3 is a tapered head tube.
  • ‘RIM_STYLE front’: [Categorical] The style of the front rim. Possible values are: ‘DISC’, ‘SPOKED’, ‘TRISPOKE’. Despite the name, trispoke class implies composite spokes but does not necessarily imply three composite spokes.
  • ‘RIM_STYLE rear’: [Categorical] The style of the rear rim. Possible values are: ‘DISC’, ‘SPOKED’, ‘TRISPOKE’. Despite the name, trispoke class implies composite spokes but does not necessarily imply three composite spokes.
  • ‘Handlebar style’: [Categorical] The style of the handlebars. Possible values are: ‘0’, ‘1’, ‘2’. 0 is a drop bar, 1 is a mountain bike bar, 2 is a bullhorn bar.
  • ‘Stem kind’: [Categorical] The style of stem. Possible values are: ‘0’, ‘1’, ‘2’. 0 is a stem that features a sharp and immediate angle away from the head tube. 1 is a stem that features a sharp angle some distance away from the head tube. 2 is a stem that features a gradual angle away from the head tube after initially extending in line with the head tube.
  • ‘Fork type’: [Categorical] The style of fork. Possible values are: ‘0’, ‘1’, ‘2’. 0 is a standard fork, 1 is a fork with mountain bike shocks, 2 is a time trial bike fork.
  • ‘Seat tube type’: [Categorical] The style of seat tube. Possible values are: ‘0’, ‘1’, ‘2’. 0 is aerodynamic, while 1 and 2 are standard round tubes with no distinction in this representation scheme.
A.6 Ethics and Societal Impact
Bike-Bench aims to advance the capabilities of Generative AI models for engineering design. In general, AI has many positives and negatives, most of which are not particularly pertinent to this work. However, we would like to acknowledge some of the pros and cons of generative models for engineering design. A principal risk of AI in engineering, particularly generative models, is safety. Generative models are usually probabilistic and when human safety is in question, even small chances of failure are unacceptable. Designs and design decisions created by AI must be held to the same (or higher) scrutiny and engineering standards as human designs. Engineering design AI also stands poised to deliver notable societal benefits. By lowering the barrier of entry to design, it may democratize design, allowing more, and especially underprivileged individuals, to participate in the design process. By increasing design throughput, it may realize better products, more efficient design, and greater design customization for individuals. We advocate for the careful and ethical use of AI in engineering design and beyond.

### v2

@@ B.4 Summary of Constraints and Objectives
The following is the description of the constraints and objectives given to the LLM, which is also a helpful human-interpretable reference:
Constraint and Objective Description Prompt: “Descriptions of all objectives and constraints in the standard BikeBench multi-objective engineering design benchmark. Some evaluation criteria are contingent on only the bike design, while others are also contingent on the conditional information. By convention, objectives are minimized at 0, with lower values being better. Constraints are also minimized, with 0 being the critical value. Larger positive values indicate more extreme constraint violation and larger magnitude negative values are more constraint-satisfying. In general, once a constraint is satisfied, we no longer care about further minimizing the value. However, achieving constraints is critical because designs that violate constraints are invalid. Thus, we must attempt to ensure that all constraints are simultaneously below zero. The 50 objectives and constraints are described as follows: Usability Score [Objective]: The predicted ’usability,’ as rated by a human, with 0 being the most usable and 1 being the least usable. Predicted by a regression model trained on human-collected ratings. Drag Force (N) [Objective]: The predicted drag force in N incurred by the cyclist in a 10 m/s headwind, as predicted by a regression model trained on computational fluid dynamics simulation data. Knee Angle Error (deg.) [Objective]: The difference between the minimum knee angle of the cyclist and the optimal reference range. May include a penalty term if the rider’s geometry is completely incompatible with the bike. Hip Angle Error (deg.) [Objective]: The difference between the torso-to-upper-leg angle of the cyclist and the optimal reference range. May include a penalty term if the rider’s geometry is completely incompatible with the bike. Arm Angle Error (deg.) [Objective]: The difference between the torso-to-arm angle of the cyclist and the optimal reference range. May include a penalty term if the rider’s geometry is completely incompatible with the bike. Arm Too Long for Bike [Constraint]: Constraint indicating that the length of the rider’s arm is longer than the length of the rider’s torso plus the distance from the saddle to the handlebars. Means the saddle is too close to the handlebars. Saddle Too Far From Handle [Constraint]: Constraint indicating that the distance from the saddle to the handlebars is longer than the length of the rider’s torso plus the length of the rider’s arm. Means the saddle is too far from the handlebars. Torso Too Long for Bike [Constraint]: Constraint indicating that the length of the rider’s torso is longer than the length of the rider’s arm plus the distance from the saddle to the handlebars. Means the saddle is too close to the handlebars. Saddle Too Far From Crank [Constraint]: Constraint indicating that the distance from the saddle to the pedals in the far position is longer than the length of the rider’s upper leg plus the length of the rider’s lower leg. Means the saddle is too far from the pedals. Upper Leg Too Long for Bike [Constraint]: Constraint indicating that the length of the rider’s upper leg is longer than the distance from the saddle to the pedals in the far position plus the length of the rider’s lower leg. Means the saddle is too close to the pedals. Lower Leg Too Long for Bike [Constraint]: Constraint indicating that the length of the rider’s lower leg is longer than the distance from the saddle to the pedals in the far position plus the length of the rider’s upper leg. Means the saddle is too close to the pedals. Cosine Distance To Embedding [Objective]: The cosine distance in the CLIP embedding space between the rendered bike image and the target text or image embedding. Mass (kg) [Objective]: The mass in kg of the bike frame, as predicted by a regression model trained on finite element analysis data. Planar Compliance Score [Objective]: A composite planar compliance score for the bike frame, as predicted by a regression model trained on finite element analysis data. Transverse Compliance Score [Objective]: A composite transverse compliance score for the bike frame, as predicted by a regression model trained on finite element analysis data. Eccentric Compliance Score [Objective]: A composite eccentric compliance score for the bike frame, as predicted by a regression model trained on finite element analysis data. Planar Safety Factor [Constraint]: Constraint quantified as 1.5 minus the safety factor under planar loading, as predicted by a regression model trained on finite element analysis data. Means the frame fails under planar loading. Eccentric Safety Factor [Constraint]: Constraint quantified as 1.5 minus the safety factor under eccentric loading, as predicted by a regression model trained on finite element analysis data. Means the frame fails under eccentric loading. Predicted Frame Validity [Constraint]: Constraint indicating some abstract issue with the frame, as predicted by a classification model trained to identify CAD models that failed to regenerate. Means the frame is invalid in an unspecified way. Saddle Height Too Small [Constraint]: Constraint indicating that the saddle height is too low to be functional. Saddle Collides With Seat Tube [Constraint]: Constraint indicating that the saddle height is so low that it collides with the top of the seat tube. Saddle Too Short [Constraint]: Constraint indicating that the length of the saddle is too short. Head Angle Over Limit [Constraint]: Constraint indicating that the head angle is over 180 degrees. Seat Angle Over Limit [Constraint]: Constraint indicating that the seat angle is over 180 degrees. Seat Post Too Short [Constraint]: Constraint indicating that the seat post doesn’t reach the seat tube given the prescribed saddle height. Seat Post Too Long [Constraint]: Constraint indicating that the seat post is so long that it hits the bottom bracket given the prescribed saddle height. Rear Wheel Inner Diameter Too Small [Constraint]: Constraint indicating that the rear wheel’s inner diameter is too small to be functional. Front Wheel Inner Diameter Too Small [Constraint]: Constraint indicating that the front wheel’s inner diameter is too small to be functional. Seat Tube Extension Longer Than Seat Tube [Constraint]: Constraint indicating that the distance from the top of the seat tube to the top tube junction is longer than the seat tube itself. Head Tube Upper Extension And Lower Extension Overlap [Constraint]: Constraint indicating that the top tube and down tubes intersect before their junctions with the head tube. Seat Stay Junction Longer Than Seat Tube [Constraint]: Constraint indicating that the distance from the top of the seat tube to the seat stay junction is longer than the seat tube itself. Non-negative Parameter Is Negative [Constraint]: Constraint indicating that at least one parameter that should be strictly positive is negative. Chain Stay Smaller Than Rear Wheel Radius [Constraint]: Constraint indicating that the chain stay length is smaller than the wheel radius, creating a collision. Chain Stay Shorter Than $B$ Drop [Constraint]: Constraint indicating that the vertical drop from the rear axle to the bottom bracket is greater than the chain stay length, creating an impossibility. Seat Stay Smaller Than Rear Wheel Radius [Constraint]: Constraint indicating that the seat stay length is smaller than the wheel radius, creating a collision. Seat Tube Collides With Rear Wheel [Constraint]: Constraint indicating that the seat tube collides with the rear wheel. Down Tube Can’t Reach Head Tube [Constraint]: Constraint indicating that the down tube isn’t long enough to reach the head tube, creating an impossibility. Rear Wheel Cutout Severs Seat Tube [Constraint]: Constraint indicating that the diameter of the aero tube cutout of the seat tube is so large such that it is completely severing the seat tube. Foot Collides With Front Wheel [Constraint]: Constraint indicating that the foot would collide with the pedal in its forward position, causing a collision when turning. Crank Hits Ground In Lowest Position [Constraint]: Constraint indicating that the crank hits the ground during its rotation. RGB Value Greater Than 255 [Constraint]: Constraint indicating that one or more frame RGB values exceed 255. Chain Stays Collide [Constraint]: Constraint indicating that the chain stays collide with each other before reaching the bottom bracket. Tube Wall Thickness Exceeds Radius [Constraint]: Constraint indicating that some tubes’ wall thickness exceeds their radius, creating an impossibility. Seat Tube Too Narrow For Seat Post [Constraint]: Constraint indicating that the seat post’s outer diameter is wider than the seat tube’s inner diameter, creating a collision. Down Tube Improperly Joins Head Tube [Constraint]: Constraint indicating that the down tube is partially disconnected from the bottom of the head tube. Top Tube Improperly Joins Head Tube [Constraint]: Constraint indicating that the top tube is partially disconnected from the top of the head tube. Top Tube Improperly Joins Seat Tube [Constraint]: Constraint indicating that the top tube is partially disconnected from the top of the seat tube. Down Tube Collides With Front Wheel [Constraint]: Constraint indicating that the down tube collides with the front wheel. Saddle Hits Top Tube [Constraint]: Constraint indicating that the saddle collides with the top tube. Saddle Hits Head Tube [Constraint]: Constraint indicating that the saddle collides with the head tube.”
B.5 Summary of Design Variables
This section contains details on BikeBench’s parametric design representation. The following is the description of the design variables given to the LLM, which is also a helpful human-interpretable reference:
Dataset Description Prompt: “Descriptions of all parameters in the standard BikeBench design representation scheme. General notes: All lengths are measured in $m$. All angles are measured in degrees. The 64 variables are described as follows: 1. ’Seatpost LENGTH’ [Continuous]: The length of the seat post. 2. ’CS textfield’ [Continuous]: The length of the chain stay tubes. 3. ’$B$ textfield’ [Continuous]: Bottom bracket drop, measured as the vertical drop from the rear axle to the center of the bottom bracket. By convention, positive values imply the bottom bracket lies below the axle. 4. ’Stack’ [Continuous]: The vertical distance from the top of the head tube to the bottom bracket. 5. ’Head angle’ [Continuous]: The angle of the head tube clockwise from horizontal, in degrees. 6. ’Head tube length textfield’ [Continuous]: The length of the head tube. 7. ’Seat stay junction0’ [Continuous]: The length along the seat tube from the top of the seat tube to the junction with the seat stays. By convention, this is measured to the center of the seat stays. 8. ’Seat tube length’ [Continuous]: The length of the seat tube. 9. ’Seat angle’ [Continuous]: The angle of the seat tube clockwise from horizontal. 10. ’DT Length’ [Continuous]: The length of the down tube. 11. ’FORK0R’ [Continuous]: Fork offset, measured as the perpendicular distance from the front axle to the head tube axis. 12. ’$B$ diameter’ [Continuous]: The diameter of the bottom bracket. 13. ’ttd’ [Continuous]: Top tube outer diameter. 14. ’dtd’ [Continuous]: Down tube outer diameter. 15. ’csd’ [Continuous]: Chain stay outer diameter. 16. ’std’ [Continuous]: Seat tube outer diameter. 17. ’htd’ [Continuous]: Head tube outer diameter. 18. ’ssd’ [Continuous]: Seat stay outer diameter. 19. ’Chain stay position on $B$’ [Continuous]: The distance along the length of the bottom bracket from its edge to the center of the chain stay tubes. 20. ’SSTopZOFFSET’ [Continuous]: The offset from the center plane of the bike of the joints connecting the seat stays to the seat tube. 21. ’MATERIAL’ [Categorical]: The material of the bike frame. Possible values are: ’ALUMINIUM’, ’STEEL’, ’TITANIUM’. 22. ’Head tube upper extension2’ [Continuous]: The length from the top of the head tube to the junction with the top tube. By convention, this is measured to the center of the top tube. 23. ’Seat tube extension2’ [Continuous]: The length from the top of the seat tube to the junction with the top tube. By convention, this is measured to the center of the top tube. 24. ’Head tube lower extension2’ [Continuous]: The length from the bottom of the head tube to the junction with the down tube. By convention, this is measured to the center of the down tube. 25. ’SEATSTAYbrdgshift’ [Continuous]: The distance along the center plane of the bike from the seat stay and seat tube junction to the seat stay bridge, if present on the bike. 26. ’CHAINSTAYbrdgshift’ [Continuous]: The distance along the center plane of the bike from the outer edge of the bottom bracket to the chain stay bridge, if present on the bike. 27. ’SEATSTAYbrdgdia1’ [Continuous]: The diameter of the seat stay bridge, if present on the bike. 28. ’CHAINSTAYbrdgdia1’ [Continuous]: The diameter of the chain stay bridge, if present on the bike. 29. ’SEATSTAYbrdgCheck’ [Boolean]: A boolean value indicating whether the seat stay bridge is present on the bike. 30. ’CHAINSTAYbrdgCheck’ [Boolean]: A boolean value indicating whether the chain stay bridge is present on the bike. 31. ’Dropout spacing’ [Continuous]: The distance between the rear dropouts. 32. ’Wall thickness Bottom Bracket’ [Continuous]: The tube wall thickness of the bottom bracket. 33. ’Wall thickness Top tube’ [Continuous]: The tube wall thickness of the top tube. 34. ’Wall thickness Head tube’ [Continuous]: The tube wall thickness of the head tube. 35. ’Wall thickness Down tube’ [Continuous]: The tube wall thickness of the down tube. 36. ’Wall thickness Chain stay’ [Continuous]: The tube wall thickness of the chain stay. 37. ’Wall thickness Seat stay’ [Continuous]: The tube wall thickness of the seat stay. 38. ’Wall thickness Seat tube’ [Continuous]: The tube wall thickness of the seat tube. 39. ’Wheel diameter front’ [Continuous]: The outer diameter of the front wheel. 40. ’RDBSD’ [Continuous]: The difference between rear wheel outer diameter and bead seat diameter, roughly approximating the tire thickness. 41. ’Wheel diameter rear’ [Continuous]: The outer diameter of the rear wheel. 42. ’FDBSD’ [Continuous]: The difference between front wheel outer diameter and bead seat diameter, roughly approximating the tire thickness. 43. ’Fork type’ [Categorical]: The style of fork. Possible values are: ’0’, ’1’, ’2’. 0 is a rigid fork, 1 is a single crown fork, 2 is a double crown fork. 44. ’Stem kind’ [Categorical]: The style of stem. Possible values are: ’0’, ’1’, ’2’. 0 is a stem that features a sharp and immediate angle away from the head tube. 1 is a stem that features a sharp angle some distance away from the head tube. 2 is a stem that features a gradual angle away from the head tube after initially extending in line with the head tube. 45. ’Handlebar style’ [Categorical]: The style of the handlebars. Possible values are: ’0’, ’1’, ’2’. 0 is a drop bar, 1 is a mountain bike bar, 2 is a bullhorn bar. 46. ’$B$ length’ [Continuous]: The length of the bottom bracket. 47. ’Wheel cut’ [Continuous]: The diameter of the cutout of the seat tube for the rear wheel, if using an aerodynamic tube type. 48. ’BELTorCHAIN’ [Boolean]: A boolean value indicating whether the bike has a chain (True) as opposed to a belt. 49. ’Number of cogs’ [Integer]: The number of cogs on the rear wheel. 50. ’Number of chainrings’ [Integer]: The number of chainrings attached to the crank. 51. ’FIRST color R_RGB’ [Continuous]: The red component of the primary paint color of the bike. 52. ’FIRST color G_RGB’ [Continuous]: The green component of the primary paint color of the bike. 53. ’FIRST color B_RGB’ [Continuous]: The blue component of the primary paint color of the bike. 54. ’RIM_STYLE front’ [Categorical]: The style of the front rim. Possible values are: ’DISC’, ’SPOKED’, ’TRISPOKE’. Despite the name, trispoke class implies composite spokes but does not necessarily imply three composite spokes. 55. ’RIM_STYLE rear’ [Categorical]: The style of the rear rim. Possible values are: ’DISC’, ’SPOKED’, ’TRISPOKE’. Despite the name, trispoke class implies composite spokes but does not necessarily imply three composite spokes. 56. ’SPOKES composite front’ [Integer]: If applicable, the number of composite spokes in the front wheel minus two (a value of 1 is a trispoke wheel). 57. ’SPOKES composite rear’ [Integer]: If applicable, the number of composite spokes in the rear wheel minus two (a value of 1 is a trispoke wheel). 58. ’SBLADEW front’ [Continuous]: If applicable, the width of the front wheel composite spokes. 59. ’SBLADEW rear’ [Continuous]: If applicable, the width of the rear wheel composite spokes. 60. ’Saddle length’ [Continuous]: The length of the saddle. 61. ’Saddle height’ [Continuous]: The vertical distance from the saddle to the bottom bracket. 62. ’Seat tube type’ [Categorical]: The style of seat tube. Possible values are: ’0’, ’1’, ’2’. 0 is aerodynamic, while 1 and 2 are standard round tubes with no distinction in this representation scheme. 63. ’Head tube type’ [Categorical]: The style of head tube. Possible values are: ’0’, ’1’, ’2’, ’3’. 0 is aerodynamic, while 1 and 2 are standard round tubes with no distinction in this representation scheme. 3 is a tapered head tube. 64. ’Down tube type’ [Categorical]: The style of down tube. Possible values are: ’0’, ’1’, ’2’. 0 is aerodynamic, while 1 and 2 are standard round tubes with no distinction in this representation scheme. ”
B.6 Conditional Tasks
Conditional information is concatenated for the LLM into a text string of the form: ‘Rider Body Dimensions: Upper leg length - [ULL], Lower leg length - [LLL], Arm length - [AL], Torso length - [TL], Neck and head length - [HNL], Torso width - [TW]. Use Case: [Road Biking/Mountain Biking/Commuting]. Bike Description: [Text Prompt]’.
An example condition from the test set is included below:
Example Test Set Condition in Text String Representation: “Rider Body Dimensions: Upper leg length - 0.3772640228, Lower leg length - 0.5058981180, Arm length - 0.6180832386, Torso length - 0.4945643246, Neck and head length - 0.3314186633, Torso width - 0.3427633643. Use Case: Road Biking. Bike Description: Track training bike with fixed gear, deep-section rims, and a stiff frame designed to transfer power effectively to the track.”
B.7 Computing Resources and Cost
All experiments were carried out on a workstation with an RTX 3090Ti GPU and a Ryzen 5900x CPU. No model training or optimization took more than 4 hours. API calls for GPT-5 queries were divided over the 100 test conditions and parallelized across 20 processes. Each LLM benchmark took approximately 1 hour to complete and cost approximately 40 USD.
B.8 Ethics and Societal Impact
BikeBench aims to advance the capabilities of generative AI models for engineering design. In general, AI has many positives and negatives, most of which are not particularly pertinent to this work. However, we would like to acknowledge some of the pros and cons of generative models for engineering design. A principal risk of AI in engineering, particularly generative models, is safety. Generative models are usually probabilistic. When human safety is in question, even small chances of failure are unacceptable. Designs and design decisions created by AI must be held to the same (or higher) engineering standards as human designs. Engineering design AI also stands poised to deliver notable societal benefits. By lowering the barrier of entry to design, it may democratize design, allowing individuals without formal training to participate in the design process. By increasing design throughput, it may realize better products, more efficient design, and greater design customization for individuals. We advocate for the careful and ethical use of AI in engineering design and beyond.